Suppose the table below shows the prices and quantities of the output produced by an economy in two years. If Year 1 is the base year, then the inflation rate in the GDP deflator is 5%.

Good a Good b

Pa Qa Pb. Qb

Year 1 $10.00 100 $4.00 500

Year 2 $9.50. 120 $4.40. 450 


1
Expert's answer
2021-02-22T10:18:10-0500
"Solution."

Here, we will have to proof if the statement that the GDP Deflator is 5% is true or false. Therefore,

Inflation rate GDP deflator

"=\frac{GDP\ deflator\ in\ year\ 2-GDP\ deflator\ in\ year\ 1}{GDP\ Deflator\ in\ year\ 1}\\
GDP\ deflator\ in\ year\ 2=\frac{Nominal\ GDP\ in\ year\ 2}{Real\ GDP\ in\ year\ 2}\\
=\frac{120\times9.50+450\times4.40}{120\times10+450\times4}\\
=\frac{3120}{3000}\times100=104\%\\
GDP\ deflator\ in\ year\ 1=\frac{Nominal\ GDP\ in\ year\ 1}{Real\ GDP\ in\ year\ 2}\\
Since, Base\ year=year\ 1\\
Nominal\ GDP=Real\ GDP\in\ year\ 1\\
Therefore, \ GDP\ deflator(year\ 1)=100\%\\
Thus, inflation\ rate\ GDP\ deflator\\
=\frac{104-100}{100}\times100\\
=4\%,"

Therefore the statement is false.